Whe Find the transfer function output for the electrical circuit below input 2H 2221F 0000 222 22 222 v(t) 232 + vo 21 VLt) 1using mesh analysis method
Answer #1

Find the transfer function output ofor the page! input electrical cruit below. 2H mo a 22 22 + Vlt) Leulu un + 2H ville) intofrom equation (i) and (ii) -276) + (4 +25)1,0) = 0 -ü) page2 (25+2) L(S) - 2/3 (5) = V(S) X 2 - 2 L(S) +(25+4) ()= 0 X (25+2)say that pages As per Electrical circuit we can Voltage across inductor is equal to VL(S) = 25 [21] ~ (S) = 25T 0.5V) s2435+1- QI(S) - 4B) + 4I. (S) + 25 [g(S) + I (S). (Page 8 21 :20 8212 vs) I, (S) I(S) 25 elle r(s) Applying KrL in mesh (61) v(s) +page solving equation 6 and equations ( theon we get (4+ 3)40 ) - (2+4) 1,60 = v) * (245) - (2+}) }(6) + (4725+) 28)=0 (012)Page 6 1 + 165² +8585² +85 I 2 (B) = [ 100 metres) (est) vi ) 1+165² +85² +85 I 2 (S) = OVO *** 10%) = (25+1) VS) [ 854161**